OpenStudy (anonymous):

Reg hex has 36 in permimeter , the apothem

OpenStudy (anonymous):

\[= 2\sqrt{3} , 3\sqrt{3} , or 6 \sqrt{}3]

OpenStudy (anonymous):

.. never mind that . Would it be 2 sqroot of 3, 3 sqroot of 3, or 6 swroot of 3?

OpenStudy (anonymous):

The apothem = length of any side/(2tan(180/number of sides)) 6/2tan(180/6)=5.196, 3sqrt3=5.196

OpenStudy (anonymous):

I'm not doing it the Tan, sine, cos way thoo . Geometry way. ._. I don't learn it like that

OpenStudy (anonymous):

Hmm I'm not sure then

OpenStudy (anonymous):

Lol Okay thanks for trying